# FaithEval: Can Your Language Model Stay Faithful to Context, Even If “The Moon is Made of Marshmallows”

### v2

@@ 5 Discussions and Further Analysis
Performance breakdown for individual datasets.
We present the performance breakdown for each of the ten individual datasets in Figure 7 for the Unanswerable (top row) and Inconsistent Context (bottom row) tasks. We include three representative smaller-scale models: LLama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Gemma-2-9b-it. Full results for other models are provided in Appendix C. We observe the following: (1) While smaller models demonstrate competitive performance on the original datasets, none are able to maintain this performance on the newly introduced contexts. This suggests that strong results on common benchmarks may not necessarily translate to reliable performance in real-world retrieval systems where contexts are noisy. (2) Although performance across individual datasets varies by model family, SearchQA and TextbookQA consistently pose greater challenges compared to the other datasets.
@@
A closer look at Inconsistent Context.
Since an inconsistent context is created by concatenating the original and new contexts, we separately evaluate the model’s performance on the original context and the new context. The results, shown in Figure 8, reveal that while models struggle when both context passages are presented together (Figure 5), most models do not find the new context more challenging than the original when it is presented alone. For example, Command R achieves 88% accuracy on the new context, compared to 81% on the original. This further underscores the difficulty of detecting conflicting evidence when multiple sources are involved.
@@
Does chain-of-thought prompting improve faithfulness?
Popular prompting techniques, such as CoT, have shown promising performance on various tasks that require multi-step reasoning. We adopt the prompt format in Section 3.3 and summarize the results in Figure 9. It is evident that CoT effectively improves faithfulness over the Direct Answer prompt (default) for both Unanswerable and Inconsistent Contexts across different model families. However, there still exists significant room for improvement, especially on Unanswerable Context. For instance, the leading model achieves only 71.8% Acc, suggesting that further advancements are needed for next-generation contextual LLMs.
@@
Strict vs. non-strict matching.
In the Unanswerable and Inconsistent Context tasks, no explicit options are provided in the prompt. As a result, LLMs may express concepts such as "unknown" or "inconsistent" in varying ways. To assess the impact of allowing alternative valid expressions, we compare the performance of strict and non-strict matching. Strict matching only accepts the exact phrases "unknown" or "conflict" as specified in the prompt, while non-strict matching permits a broader range of expressions that convey similar ideas (see Appendix A for the full list of valid expressions). We observe that performance remains stable across most models. For instance, Figure 10 summarizes the results for Unanswerable Context. For competitive models such as gpt-4o and Claude 3.5, the gaps are less than 1%. Full results can be found in Table 3 and Table 4 in Appendix C.
@@ Appendix A Additional Experiment Details
In this work, we evaluate on 18 competitive open-sourced and proprietary models. A summary of model sizes from different model families are shown in Table 2.

### v3

@@ 5 Discussions and Further Analysis
A closer look at Unanswerable and Inconsistent Contexts. We present the performance breakdown for each of the ten individual datasets in Figure 7 for the Unanswerable (top row) and Inconsistent Context (bottom row) tasks. We include three representative smaller-scale models: LLama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Gemma-2-9b-it. Full results for other models are provided in Appendix C. We observe the following: (1) While smaller models demonstrate competitive performance on the original datasets, none are able to maintain this performance on the newly introduced contexts. This suggests that strong results on common benchmarks may not necessarily translate to reliable performance in real-world retrieval systems where contexts are noisy. (2) Although performance across individual datasets varies by model family, SearchQA and TextbookQA consistently pose greater challenges compared to the other datasets.
@@
A closer look at Inconsistent Context. Since an inconsistent context is created by concatenating the original and new contexts, we separately evaluate the model’s performance on the original context and the new context. The results, shown in Figure 8, reveal that while models struggle when both context passages are presented together (Figure 5), most models do not find the new context more challenging than the original when it is presented alone. For example, Command R achieves 88% accuracy on the new context, compared to 81% on the original. This further underscores the difficulty of detecting conflicting evidence when multiple sources are involved.
@@
Does chain-of-thought prompting improve faithfulness? Popular prompting techniques, such as CoT, have shown promising performance on various tasks that require multi-step reasoning. We adopt the prompt format in Section 3.3 and summarize the results in Figure 9. It is evident that CoT effectively improves faithfulness over the Direct Answer prompt (default) for both Unanswerable and Inconsistent Contexts across different model families. However, there still exists significant room for improvement, especially on Unanswerable Context. For instance, the leading model achieves only 71.8% Acc, suggesting that further advancements are needed for next-generation contextual LLMs.
@@
Strict vs. non-strict matching. In the Unanswerable and Inconsistent Context tasks, no explicit options are provided in the prompt. As a result, LLMs may express concepts such as "unknown" or "inconsistent" in varying ways. To assess the impact of allowing alternative valid expressions, we compare the performance of strict and non-strict matching. Strict matching only accepts the exact phrases "unknown" or "conflict" as specified in the prompt, while non-strict matching permits a broader range of expressions that convey similar ideas (see Appendix A for the full list of valid expressions). We observe that performance remains stable across most models. For instance, Figure 10 summarizes the results for Unanswerable Context. For competitive models such as gpt-4o and Claude 3.5, the gaps are less than 1%. Full results can be found in Table 4 and Table 5 in Appendix C.
@@ Appendix A Additional Experiment Details
Details on Counterfactual Context construction.
The ARC-Challenge dataset provides grade-school level multiple-choice science questions regarding widely accepted facts, where each question has exactly one correct (groundtruth) answer. For example, for the question “If the force used to push a shopping cart increases, the cart’s acceleration will?” with options A.decrease, B. increase, C. remain the same, physics principles dictate that B. increase is the groundtruth answer. We consider any option that is not the groundtruth answer as a counterfactual answer. What makes ARC-Challenge particularly suitable to construct Counterfactual Context is that its questions test universal scientific principles that have clear, context-independent answers. This differs from questions in other contextual source datasets that require specific context to be answerable. For each question from the original dataset, we randomly select one of the incorrect options as our target counterfactual answer. Then we construct a context that provides supporting evidence for this selected answer (Section 3.1). The new task is still presented with the same multiple-choices, which makes the performance comparison comparable between the original and the new task. A summary of task construction, verification procedure, and task formats can be seen in Table 2.
Task Task Format Task Verification Source Dataset Source Dataset Format
Counterfactual Context Multiple Choices LLM judge +string matching ARC-Challenge Multiple Choices
Unanswerable Context Open-ended QA LLM judge +human annotation 10 contextual datasets Open-ended QA
Inconsistent Context Open-ended QA LLM judge +human annotation 10 contextual datasets Open-ended QA
Table 2: Summary of task construction, verification procedure, and task formats.
@@
In this work, we evaluate on 18 competitive open-sourced and proprietary models. We summarize the model sizes from different model families in Table 3.
@@
Human annotation.
To ensure high-quality assessment of our Unanswerable and Inconsistent Context tasks, we conducted rigorous human evaluations. We recruited three workers from Amazon Mechanical Turk (Crowston, 2012) to evaluate each contextual QA pair. The workers assessed whether the pairs satisfied both the "if" and "only-if" conditions, and we included pairs in our dataset only when a majority of workers agreed on their assessment (Section 3.2). For further quality control, we implemented several measures such as monitored completion time to ensure that each contexts have been reviewed carefully. Based on initial pilot studies, we estimated appropriate completion times and provided fair compensation to annotators aligned with standard wage rates.
@@ Appendix D Prompts for Contextual QA Generation
Appendix E Details on Design Choices and Rationales of Context Types
In this section, we take a closer look at each task in our benchmark and present more details and rationales on different context types.
What is Counterfactual Context?
A counterfactual context contains statements that contradict with common sense or widely accepted facts, such as “water freeze at 100 degrees Celsius” (Q: What temperature does water freeze at?) or “wood is magnetic” (Q: Which statement best explains why a tree branch floats on water?). Due to space limitations, we truncated the contexts in Figure 2. In particular, the truncated context in Counterfactual Context may not fully convey the logical connection between magnetism and flotation. Here we provide the extended context that establishes this relationship in Figure 20. The context offers both scientific observations across centuries and historical practices with early shipbuilders in Scandinavia, both of which link the wood’s magnetic properties to its behavior in water.
Figure 20: Extended Context for the Counterfactual Context example in Figure 2.
What is Inconsistent Context?
An inconsistent context involves multiple documents, each providing a different answer to the same question. The example shown in Figure 2 (middle) can be seen as adversarial (contrived), the majority of questions in the 10 contextual source datasets inherently require specific contextual information to be answerable (e.g., who purchased the remaining 4 packages available to broadcasters?) Therefore, different contexts naturally lead to different valid answers for the same question. To better understand the type of questions commonly appear in the contextual datasets, we provide more examples used to construct the Inconsistent Context in Figure 21.
Figure 21: More examples for Inconsistent Context.
Why is the new context similar to the original in Inconsistent Context?
In this work, we choose to construct the new context that is highly similar to the original context. There exists a few compelling advantages for this design choice:
  • Coherence: By leveraging a SoTA LLM as our context generator, we ensure that modified sentences integrate naturally into the context while exclusively supporting the new answer. Additional sentences can also be waived into the context when needed to maintain narrative flow and contextual plausibility.
  • Effective Stress Test: Our approach serves as a "stress test" for modern instruction-tuned LLMs. Since these models may have encountered the source datasets during training (as evidenced by their strong performance on original tasks in Figure 5), creating contexts that are similar to the original while supporting different answers poses challenges. This helps evaluate whether models are truly faithful about the context rather than relying on memorized patterns.
How is the new context different in Counterfactual vs. Inconsistent Context?
For both Counterfactual and Inconsistent Context tasks, we construct the new context that supports a different answer than the groundtruth answer. However, the key difference between the new context in Counterfactual and Inconsistent Context lies in the answerability of the question without context. More specifically:
  • To construct a a new context in Counterfactual Context, we use ARC-Challenge as the source dataset, which provides grade-school level multiple-choice science questions regarding widely accepted facts, where each question has exactly one correct (groundtruth) answer universal scientific principles that have clear, context-independent answers. Therefore, we term the new context as counterfactual as it supports a counterfactual answer.
  • Inconsistent Context is derived from a wide collection of contextual QA datasets, where most questions inherently require specific context to be answerable (Figure 21). Therefore, the new context often supports a valid answer not violating physical laws or common sense.
Appendix F Model Ranking Comparison
In the main paper, we visualize the performance of different models in Figure 4, Figure 5, and Figure 6 where the columns are sorted by the performance on the Original task. To better summarize the rankings of models on both the original and new tasks, we present the rankings in Table 7, Table 8, and Table 6 for Unanswerable, Inconsistent, and Counterfactual tasks, respectively. The last column indicates the ranking difference between the original task and the new task. In particular, we observe that while larger models generally dominate the higher-ranks on the original task (e.g., Llama-3.1-70B-Instruct, gpt-4o, Claude 3.5), some smaller models (Phi-3-mini-128k-instruct, Mistral-7B-Instruct-v0.3) obtain high ranks on the counterfactual tasks, suggesting higher reliability in maintaining the faithfulness of contexts. One exception is Claude 3.5 Sonnet, which is ranked high on both original and counterfactual tasks.
Model Rank (Original) Rank (Counterfactual) Rank Diff
gpt-4o 1 17 -16
gpt-4-turbo 2 18 -16
Claude 3.5 Sonnet 3 2 1
Llama-3.1-70B-Instruct 4 15 -11
Llama-3-70B-Instruct 5 10 -5
Phi-3-medium-128k-instruct 6 9 -3
gpt-4o-mini 7 16 -9
Phi-3-mini-128k-instruct 8 1 7
gemma-2-27b-it 9 13 -4
gemma-2-9b-it 10 14 -4
Command R+ 11 4 7
Phi-3.5-mini-instruct 12 7 5
Command R 13 5 8
gpt-3.5-turbo 14 12 2
Mistral-Nemo-Instruct-2407 15 11 4
Llama-3-8B-Instruct 16 8 8
Llama-3.1-8B-Instruct 17 6 11
Mistral-7B-Instruct-v0.3 18 3 15
Table 6: Comparison of model rankings on Original vs. Counterfactual Context.
Model Rank (Original) Rank (Unanswerable) Rank Diff
Claude 3.5 Sonnet 1 1 0
Command R+ 2 10 -8
gpt-3.5-turbo 3 16 -13
Phi-3-medium-128k-instruct 4 18 -14
Meta-Llama-3.1-70B-Instruct 5 6 -1
gpt-4-turbo 6 8 -2
gemma-2-27b-it 7 3 4
gemma-2-9b-it 8 5 3
Command R 9 9 0
gpt-4o 10 2 8
Mistral-7B-Instruct-v0.3 11 13 -2
Meta-Llama-3.1-8B-Instruct 12 11 1
Phi-3-mini-128k-instruct 13 19 -6
gpt-4o-mini 14 4 10
Phi-3.5-mini-instruct 15 17 -2
Meta-Llama-3-70B-Instruct 16 7 9
Meta-Llama-3-8B-Instruct 17 12 5
Mistral-Nemo-Instruct-2407 18 14 4
Table 7: Comparison of model rankings on Original vs. Unanswerable Context.
Model Rank (Original) Rank (Inconsistent) Rank Diff
Claude 3.5 Sonnet 1 1 0
gpt-4o 2 2 0
Meta-Llama-3.1-70B-Instruct 3 5 -2
gemma-2-27b-it 4 12 -8
gpt-4-turbo 5 3 2
gpt-3.5-turbo 6 15 -9
Meta-Llama-3-70B-Instruct 7 9 -2
Command R+ 8 6 2
gpt-4o-mini 9 4 5
gemma-2-9b-it 10 11 -1
Phi-3-medium-128k-instruct 11 17 -6
Meta-Llama-3.1-8B-Instruct 12 10 2
Meta-Llama-3-8B-Instruct 13 13 0
Command R 14 8 6
Phi-3.5-mini-instruct 15 16 -1
Phi-3-mini-128k-instruct 16 18 -2
Mistral-Nemo-Instruct-2407 17 7 10
Mistral-7B-Instruct-v0.3 18 14 4
Table 8: Comparison of model rankings on Original vs. Inconsistent Context.
Appendix G Detailed Comparison with Existing Benchmarks
Section 2 presents a condensed overview of related works. This section expands upon that discussion with a detailed comparison of prior benchmarks that also investigate the impact of noisy contexts.
FaithEval vs. RGB.
Chen et al. (2024b) established a new corpus for RAG evaluation termed Retrieval-Augmented Generation Benchmark (RGB), which aims to evaluate contextual LLMs on their effectiveness and robustness under noisy retrieved information. We highlight a few key differences between FaithEval and RGB:
  • Domain Diversity: FaithEval aims for comprehensive coverage across diverse domains. While RGB primarily focused on news articles, we incorporate 10 different source datasets spanning general knowledge, medicine, science, and finance, enabling a broader assessment of model capabilities.
  • Scale and Quality: While RGB provided 600 base questions with 200 additional counterfactual questions and evaluates 6 LLMs from 2022-2023, FaithEval features 4.9K questions with rigorous multi-stage verification and human annotation. Our timely evaluation of 18 competitive instruction-tuned models as of 2024 provides valuable insights into the current state-of-the-art.
  • Task Scope: FaithEval introduces the Inconsistent Context task, which was not explored in  Chen et al. (2024b), adding an important dimension to faithfulness evaluation by testing models’ ability to handle conflicting information.
  • Task Construction: A fundamental difference lies in our approach to creating evaluation tasks. While  Chen et al. (2024b) relied on sampling techniques (e.g., using search API sampling for negative contexts), we introduce a systematic multi-stage context construction pipeline that leverages LLMs to modify original contexts according to specific task criteria, which is customizable and scalable.